OpenStudy (anonymous):

how do you simplify (2^-2 + 5^-1)^-2

OpenStudy (anonymous):

okay. what does 2^-2 mean? do you know?

OpenStudy (anonymous):

well? do you need my help? because I need your feedback to help you, erin.

OpenStudy (amistre64):

you write it in arabic :)

OpenStudy (anonymous):

Alright, the problem is the following:\[\frac{1}{(2^{-2} + 5^{-1})^2}\] when you have a negative power, you can make it positive by simply changing its position ^_^. So now you have : \[\frac{1}{\frac{1}{2^2} + \frac{1}{5}}= \frac{1}{\frac{1}{4} + \frac{1}{5}}\] Now compute the denominator, then work it out as a simple fraction : \[=\frac{1}{\frac{5+4}{20}}=\frac{1}{\frac{9}{20}} = \frac{20}{9}\] I hope this made the concept of the idea clearer ^_^

OpenStudy (anonymous):

lol wait, I forgot the square part, but you'll figure it out :)sorry ^_^